2 points H is pointing 220 relative to the VERTICAL in quadrant 3. Hy 89. Determine the magnitude and x-component of H. 2.+ -132 points gnitude of 29 and is located at 2000 on the force table. Determine the magnitude & angle (between o & 3609) of -A. -A (magnitude) @ o (angle) Rx 93 & Ry 22. Determine the magnitude and angle of R. 2 Direction of R o relative to the VERTICAL in quadrantChoose quadrant
Answer #1

1 )

H angle is 22o

Hy = 89

here given

\theta = tan-1 ( Hy / Hx )

22 = tan-1 ( Hy / Hx )

22 = tan-1 ( 89 / Hx )

tan 22 = 89 / Hx

Hx = 89 / tan 22

Hx = 220.2827

then H = ( Hx2 + Hy2 )1/2

= ( 220.28272 + 892 )1/2

H = 237.5825